OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84.

OpenStudy (anonymous):

since, y varies directly with x. therefore \[y \alpha x \rightarrow y=kx...(k=constant)\] Now at y = -7 when x = -1, k=? \[-7=ktimes (-1).\rightarrow k= \frac{-7}{-1} =7\] thus our eq of cvariation becomes y= 7x Now at y=-84, x=? -84=7x x=-84/7 x=-12

OpenStudy (anonymous):

@francecsamusco